On August 1, John issued a $600,000, semi-annual, 5 year, 3.5% bond. The market rate for similar bonds on that day was 4.0%. John uses the effective interest method to record the amortization of discounts. He has decided to report net bonds on the balance sheet, instead of reporting the bond and its discount separately.

What does net bonds look like on the balance sheet and how do I get it from bonds payable and discounts?
Answer #1
Total interest on issued bond @ 3.5% 600000*10*3.5% = 210000
Total interest on similar bond @ 4% 600000*10*4% = 240000
Discount on issue 210000-240000 = 30000
So net bond on book will be 600000-30000 = 570000
we have to deduct the discount from bond value to get the net bond value